# A Survey on Brain-inspired Deep Learning via Predictive Coding

### v2

@@ Abstract
Artificial intelligence (AI) is rapidly becoming one of the key technologies of this century. The majority of results in AI thus far have been achieved using deep neural networks trained with the error backpropagation learning algorithm. However, such an algorithm has always been considered biologically implausible. To this end, recent works have studied learning algorithms for deep neural networks inspired by the neurosciences. One such theory, called predictive coding (PC), has shown promising properties that make it potentially valuable for the machine learning community: it can model information processing in different areas of the brain, can be used in cognitive control and robotics, has a solid mathematical foundation in variational inference, and performs its computations asynchronously. Inspired by such properties, works that propose novel PC-like algorithms are starting to be present in multiple sub-fields of machine learning and artificial intelligence at large. Here, we survey such efforts by first providing a broad overview of the history of PC to provide a common ground for the understanding of the recent developments, then by surveying current efforts and results, and concluding with a large discussion of possible implications and ways forward.
@@ 1 Introduction
The machine learning community is developing and producing models that push the boundaries of the field on a weekly basis: we have witnessed considerable breakthroughs in the fields of generative artificial intelligence (AI) [1, 2, 3, 4], game playing [5, 6], and text-generation [7, 8, 9]. These results reflect over a decade of advances in the field, made possible by the joint effort of tens of thousands of researchers and engineers who have built on seminal work [10, 11] in order to improve the performance of deep artificial neural networks trained with the error backpropagation algorithm [12, 13]. It may thus come as a surprise that research on alternative training methods for machine learning is more active than ever, with many works often co-authored by the same scientists who pioneered backpropagation-based schemes [14, 15, 16, 17]. These lines of research, however, are not contradictory but complementary: the increasingly impressive results of standard deep neural networks have also highlighted several significant limitations, that may be addressed via alternative training methods. There is, in fact, a common belief that, to accelerate progress in AI, we must invest in fundamental research, inspired by the findings and ideas in computational neuroscience [18, 17, 19]. To this end, in this review, we focus on a specific approach, inspired from a theory of learning and perception in the brain called predictive coding (PC) [20, 21, 22]. Models of PC exhibit several valuable properties, the main one being the locality of the operations, meaning that every update of the parameters is performed using only the information of adjacent neurons or synapses in an Hebbian-way [23]. Differently from backprop, that is a sequential algorithm, locality allows for a full parallelization of the operations, making it potentially more suitable for applications in neuromorphic computing. Locality also allows the training of artificial neural networks with any given topology [24, 25]. This facilitates the training of models with arbitrarily entangled structures, similar to the biological networks that constitute our brains and underwrite our level of intelligence [26]. Furthermore, PC networks are naturally more robust in relation to standard models trained with backprop, due to the relaxation process that spreads error signals across the whole network [27]. In supervised learning, this process emulates implicit gradient descent [28], is naturally more stable, and is less reliant on hyperparameter searches. This allows predictive-coding-based models to perform better in tasks that are more likely to be faced by natural agents, while regularizing the underlying energy landscape [29] and avoiding troublesome phenomena, such as vanishing and exploding gradients [27, 30].
Structure of the Survey.
This article is organized as follows. In Section 2, we present the general theoretical characterization of PC as (the inversion of) a generative model that minimizes variational free energy, provide an hystorical introduction to the topic, as well as an informal definition of what characterizes a predictive coding algorithm in the literature. Section 3 is quite technical, as we dive into the mathematical formulations to study and dissect three key computational frameworks that have emerged over the past three decades: the classical setup proposed by Rao and Ballard, neural generative coding, and biased competition with divisive modulation. Following this, in Section 4, we review the machine learning problems that PC has been tested on, such as supervised learning, natural language processing, computer vision, temporal and continual learning, and reinforcement learning and control for robotics and cognitive architecture design. In Section 5, we examine key neuroscience studies of PC, and, in Sections 6 and 7, we consider available software tools and libraries that support PC, and discuss how research on hardware could realize the full potential of PC or similar algorithms. To conclude, in Section 8, we discuss important future directions, as well as the grand challenges that lie ahead.
@@ 2 Generative Models and Predictive Coding
Given an observation $\mathbf{o}$, and a latent variable $\mathbf{x}$, a generative model is described by the marginal probability $p(\mathbf{o},\mathbf{x})=p(\mathbf{o}\mid\mathbf{x})p(\mathbf{x})$. In the spirit of one of Richard Feynman’s famous, oft-quoted statements, What I cannot create, I do not understand, this family of models aims to understand patterns found within a dataset by learning how the dataset can be generated. Making these types of models work for specific tasks entails two steps: (1) building the correct structure of the generative model, that requires making assumptions on the structure of the function that generates plausible datapoints using latent variables as inputs, and (2) inverting the ensuing generative model, that requires understanding what is the best or optimal inference (i.e., belief) updating scheme that enables us to invert such a probabilistic generative model.
To invert a generative model means to map from consequences (i.e., content or data) to causes (i.e., latent states), that is, to identify the posterior $p(\mathbf{x}\mid\mathbf{o})$. After the advent of variational autoencoders (VAEs) [31], the inversion behind most, if not all, modern, successful generative AI models is amortized, meaning that the computation of the (approximate) posterior is performed via a forward pass on a specifically trained deep neural network, called the encoder, as sketched in Fig. 1 (left). However, statistical models have been around for many decades, and provided many techniques that facilitate the inversion of a generative model that does not rely on an amortized scheme, that is, can be defined as sketched in Fig. 1 (right). We now describe how a generative model based on PC first performs inference – that describes how the latent variables are updated – and learning – that describes how the parameters of the model are updated.
Inference.
Given an observation $\mathbf{o}$, the goal here is to approximate a posterior distribution $p(\mathbf{x}\mid\mathbf{o})$ using the joint distribution of observations and latent variables $\mathbf{x}$. As $\mathbf{x}$ is often of lower dimension than $\mathbf{o}$, this can be read as a compression mechanism, where compression can be defined as the process of assigning a lower dimensional code to a particular observation. This coding mechanism is what originally gives the name to the PC framework. In fact, the first algorithmic formulation of PC dates back to the $1950$’s, and was a compression algorithm for time series data [32, 33, 34]. The first instances of PC in neuroscience also treated it as a pure inference algorithm, with no learning involved, as the goal was to model how spatial information was encoded in the retina [35].
Learning.
A major paradigm shift, however, happened several years later, with the work of Rao and Ballard, which employed PC as a learning algorithm, instead of just an inference one to model hierarchical processing in the visual cortex [20]. In this formulation, synaptic weights are updated after inference to minimize the prediction error of the model. The connection with variational inference was discovered a couple of years later, and connected PC to hierarchical Gaussian generative models [36, 37]. As a result of these independent, but deeply related lines of work, a recent general theory has unified these views by showing that these iterative updating schemes can be described as a process of minimizing a variational free energy [38], also known in machine learning as an evidence bound [39]. A sketch of a timeline of the main breakthroughs in the history of PC is given in Fig. 2.
@@
The ability to memorize a training set, learn its patterns, and generalize to an unseen test set is the primary goal of any machine learning algorithm. The variational free energy is a function that encodes the two factors that model this behavior, one that encourages the fitting of the training dataset (training accuracy), and one that penalizes the complexity of the generative model, facilitating a better performance on unseen data. This enforces regularity, as it facilitates convergence towards the minimally complex model that provides an accurate fit of a specific dataset, aligned with Occam’s razor formulations [40]. As an example, let us assume that we have a discrete generative model, expressed as the probability $p(\mathbf{o},\mathbf{x})=p(\mathbf{o}\mid\mathbf{x})p(\mathbf{x})=p(\mathbf{x}\mid\mathbf{o})p(\mathbf{o})$ , where $\mathbf{x}$ denotes a latent vector, and $\mathbf{o}$ is an observed data point. Given an observation $\mathbf{o}$, and an approximate posterior distribution $q_{\theta}(\mathbf{x})$, parameterized by a vector $\theta$, the (negative) variational free energy is defined as follows:
@@
$E(p,q,\mathbf{o})=\underbrace{\sum_{\mathbf{x}}q_{\theta}(\mathbf{x})\text{log}\bigg{(}\frac{q(\mathbf{x})}{p(\mathbf{x})}\bigg{)}}_{\text{Complexity}}\ +\ \underbrace{\sum_{\mathbf{x}}q_{\theta}(\mathbf{x})\text{log}\bigg{(}\frac{1}{p(\mathbf{o}\mid\mathbf{x})}\bigg{)}}_{\text{Accuracy}}\,,$
A typical schema to minimize such a variational free energy is the expectation-maximization (EM) algorithm, where inference is first used to find the latent vector that minimize $E(p,q,\mathbf{o})$, and then the parameters $\theta$ are further updated to follow the same goal.
@@
2.1 What Defines Predictive Coding?
In the following section, we will discuss in detail different variations of PC proposed in the recent literature. However, when can an algorithm be defined as PC? Here, we address this question by providing an informal definition of what predictive coding is today. Let $\mathbf{o}=g(\mathbf{x},\mathbf{W})$ be a generative model, where a vector of latent variables $\mathbf{x}$ and a set of parameters $\mathbf{W}$ are used to generate an observation $\mathbf{o}$. In our case, we consider generative models with $L$ layers, whose marginal probability has the following dependencies:
@@ 2.1 What Defines Predictive Coding?
We further consider the probability distribution $p(\mathbf{x}_{\ell}\mid\mathbf{x}_{\ell+1})$ to be a multivariate Gaussian distribution, whose mean is given by a transformation $g_{\ell}$ of the latent variables of the level above. In the majority of the literature, the map $g_{\ell}$ is a composition of an activation function (such as ReLU), and a transformation matrix $\mathbf{W}_{\ell}$ (with bias $\mathbf{b}_{\ell}$ possibly folded into this matrix), which results in a linear map. As a result, we then, more formally, arrive at the following:
$p(\mathbf{x}_{L})=\mathcal{N}(\mathbf{x}_{L},\mathbf{\Sigma}_{L}),\ \ \ \ p(\mathbf{x}_{\ell}\mid\mathbf{x}_{\ell+1})=\mathcal{N}(g_{\ell+1}(\mathbf{x}_{\ell+1}),\mathbf{\Sigma}_{\ell}).$ (2)
@@
Given the above, we now have the structure of the main object of study that centrally characterizes the PC literature, that is, a hierarchical generative model, which lives in a continuous state space and whose probability distributions are Gaussian in nature. However, what also defines PC is the process that is used to invert the generative model, which means estimating $p(\mathbf{x}_{1},\dots,\mathbf{x}_{L}\mid\mathbf{o})$ via an approximate posterior $q(\mathbf{x}_{1},\dots,\mathbf{x}_{L})$. To do this, we need to appeal to two different approximation results: first, we leverage a mean-field approximation, which allows us to assume that the variational posterior factorizes into conditionally independent posteriors $q(\mathbf{x}_{\ell})$; then, through the Laplace approximation, we approximate posterior distributions as Gaussian forms [44]. At this point, it is now possible to minimize the resulting variational free energy. In essence, what results is an inversion mechanism that has the same dynamics as the original computational model in Rao and Ballard [20], which was notably developed before PC was cast as variational learning and inference. The minimization of variational free energy can either use gradient descent or fixed point iterations [45]. The latter follows from the fact that Gaussian priors are conjugate to Gaussian posteriors. In either case, the gradients of variational free energy are just prediction errors. This means, to evaluate requisite free energy gradients one has to evaluate prediction errors. One might argue that PC is the explicit (or implicit) use of prediction errors for the inversion of generative models (under Gaussian assumptions about random effects).
We have now introduced all of the core concepts needed to provide a definition of PC. This definition could be used as an umbrella for all of the variations on PC in the literature.
@@
    it maximizes the model evidence $p(\mathbf{o})$ by minimizing a variational free energy,
@@
    the posterior distributions of the nodes of the heterarchical structure are factorized via a mean-field approximation, and
@@
Note that the above definition does not say anything explicitly about prediction error or properties such as locality, which, as mentioned earlier, are commonly used to describe PC. This is because they are not foundational to PC but rather consequences of the commitment to the aforementioned generative model: the mean field approximation enforces independence, and hence, results in locality in the update rules; the Laplace assumption lends the consequent variational free energy a quadratic form, meaning that its gradients are linear prediction errors. It is crucial to note that the above definition of PC does not mean, however, that the algorithms surveyed here are optimal and cannot be improved: the above definition is quite general and does not impose any constraint on the exact computation of the posteriors as well as the optimization technique(s) used to minimize the variational free energy. The goal of this work is to survey existing PC methods and to anticipate future research in the field.
3 Popular Implementations of Predictive Coding

### v3

# Brain-inspired Computational Intelligence via Predictive Coding

@@ Abstract
Artificial intelligence (AI) is rapidly becoming one of the key technologies of this century. The majority of results in AI thus far have been achieved using deep neural networks trained with a learning algorithm called error backpropagation, always considered biologically implausible. To this end, recent works have studied learning algorithms for deep neural networks inspired by the neurosciences. One such theory, called predictive coding (PC), has shown promising properties that make it potentially valuable for the machine learning community: it can model information processing in different areas of the brain, can be used in control and robotics, has a solid mathematical foundation in variational inference, and performs its computations asynchronously. Inspired by such properties, works that propose novel PC-like algorithms are starting to be present in multiple sub-fields of machine learning and AI at large. Here, we survey such efforts by first providing a broad overview of the history of PC to provide common ground for the understanding of the recent developments, then by describing current efforts and results, and concluding with a large discussion of possible implications and ways forward.
@@ 1 Introduction
The machine learning community is developing and producing models that push the boundaries of the field on a weekly basis: we have witnessed considerable breakthroughs in the fields of generative artificial intelligence (AI) [1, 2, 3, 4], game playing [5, 6], and text-generation [7, 8, 9]. These results reflect over a decade of advances in the field, made possible by the joint effort of tens of thousands of researchers and engineers who have built on seminal work [10, 11] in order to improve the performance of deep artificial neural networks trained with the error backpropagation algorithm [12, 13]. It may thus come as a surprise that research on alternative training methods for machine learning is more active than ever, with many works often co-authored by the same scientists who pioneered backpropagation-based schemes [14, 15, 16, 17]. These lines of research, however, are not contradictory but complementary: the increasingly impressive results of standard deep neural networks have also highlighted several significant limitations, that may be addressed via alternative training methods. There is, in fact, a common belief that, to accelerate progress in AI, we must invest in fundamental research, inspired by the findings and ideas in computational neuroscience [18, 17, 19]. To this end, in this review, we focus on a specific approach, inspired by a theory of learning and perception in the brain called predictive coding (PC) [20, 21, 22]. Models of PC exhibit several valuable properties, the main one being the locality of the operations, meaning that every update of the parameters is performed using only the information of adjacent neurons or synapses in a Hebbian way [23]. Differently from backprop, which is a sequential algorithm, locality allows for a full parallelization of the operations, making it potentially more suitable for applications in neuromorphic computing. Locality also allows the training of artificial neural networks with any given topology [24, 25]. This facilitates the training of models with arbitrarily entangled structures, similar to the biological networks that constitute our brains and underwrite our level of intelligence [26]. Furthermore, PC networks have shown to be robust and stable due to the relaxation process that spreads error signals across the whole network, and emulates implicit gradient descent [27, 28, 29], allowing them to perform well in tasks that are more likely to be faced by natural agents.
@@ 2 Generative Models and Predictive Coding
Generative modeling provides a principled statistical framework for understanding complex data by decomposing observations into latent causes and observable effects. At its foundation, a generative model specifies a joint probability distribution over observed data $\mathbf{o}$ and latent variables $\mathbf{x}$, formally expressed as $p(\mathbf{o},\mathbf{x})=p(\mathbf{o}\mid\mathbf{x})p(\mathbf{x})$. This factorization separates the generative process—how latent causes produce observations through the likelihood $p(\mathbf{o}\mid\mathbf{x})$—from prior beliefs about the structure of those causes, encoded in $p(\mathbf{x})$. To build intuition, consider modeling handwritten digits: the latent variables $\mathbf{x}$ might capture high-level concepts like digit identity, writing style, or stroke thickness, while the likelihood $p(\mathbf{o}\mid\mathbf{x})$ describes how these abstract concepts combine to produce the pixel intensities we observe in an image. The prior $p(\mathbf{x})$ encodes beliefs about which combinations of these concepts are more or less likely to occur naturally.
The power of generative modeling lies in its ability to address two fundamental computational problems. The first is inference: given an observation $\mathbf{o}$, what are the most plausible latent causes that generated it? This requires computing the posterior distribution $p(\mathbf{x}\mid\mathbf{o})$ using Bayes’ rule. The second is learning: given a dataset of observations, how should we adjust the model parameters to better capture the underlying data distribution? This typically involves maximizing the model evidence $\log p(\mathbf{o})$ with respect to the model parameters, integrating over all possible latent variable configurations. In practice, both inference and learning present significant computational challenges: computing the posterior $p(\mathbf{x}\mid\mathbf{o})$ requires evaluating the often intractable marginal likelihood $p(\mathbf{o})$, while learning requires optimizing the model evidence $\log p(\mathbf{o})$ over parameter space. These computational bottlenecks have motivated the development of various approximation schemes, each making different trade-offs between accuracy and computational efficiency.
Variational Free-Energy.
@@
The ability to memorize a training set, learn its patterns, and generalize to an unseen test set is the primary goal of any machine learning algorithm. The variational free energy provides a tractable objective that balances two critical factors: one that encourages fitting the training dataset (training accuracy), and one that penalizes the complexity of the inferred latent variables, facilitating better performance on unseen data. This enforces regularity, as it facilitates convergence towards the minimally complex model that provides an accurate fit of a specific dataset, aligned with Occam’s razor formulations [31]. Given an observation $\mathbf{o}$ and an approximate posterior distribution $q_{\theta}(\mathbf{x})$ parameterized by $\theta$, the evidence lower bound (ELBO) or negative variational free energy is defined as:
@@
$\mathcal{L}(\theta,\mathbf{o})=\underbrace{\mathbb{E}_{q_{\theta}(\mathbf{x})}[\log p(\mathbf{o}\mid\mathbf{x})]}_{\text{Accuracy}}-\underbrace{D_{KL}[q_{\theta}(\mathbf{x})\|p(\mathbf{x})]}_{\text{Complexity}}$
This objective provides a lower bound on the log marginal likelihood: $\log p(\mathbf{o})\geq\mathcal{L}(\theta,\mathbf{o})$. The accuracy term encourages the model to assign high likelihood to the observed data under the inferred latent variables, while the complexity term penalizes deviations of the approximate posterior from the prior, preventing overfitting and ensuring generalization. A typical schema to maximize this bound is the expectation-maximization (EM) algorithm, where inference first finds the variational parameters $\theta$ that maximize $\mathcal{L}(\theta,\mathbf{o})$, and then the generative model parameters are updated to follow the same goal.
Approximate Posterior.
Today’s most prominent approach, the variational autoencoder (VAE) [30], uses amortized inference to compute the approximated posterior. This means that $q_{\theta}(\mathbf{x})$ is the output of the forward pass of a specifically trained neural network encoder, as sketched in Fig. 1 (left). However, an alternative class of methods performs inference iteratively, without relying on a separate encoder network. Predictive coding (PC) represents one such approach, using iterative optimization to approximate the posterior while maintaining the same generative model throughout both inference and learning, as sketched in Fig. 1 (right). This computational framework, while having roots in signal processing and neuroscience, provides a principled approach to generative modeling that we now describe in more detail.
2.1 What Defines Predictive Coding?
In the PC framework, given an observation $\mathbf{o}$, the goal is to approximate the posterior distribution $p(\mathbf{x}\mid\mathbf{o})$ through iterative optimization rather than amortized inference. Historically, this process was developed as a compression mechanism, where compression is defined as the process of assigning a lower-dimensional latent code to a particular observation. The first algorithmic formulation of PC dates back to the 1950s as a compression algorithm for time series data [32, 33, 34], and it is this coding that gave the framework its name. The first instances of PC in neuroscience also treated it as a pure inference algorithm, with no learning involved, as the goal was to model how spatial information was encoded in the retina [35].
A major paradigm shift occurred several years later with the work of Rao and Ballard, which employed PC as a learning algorithm in addition to inference, modeling hierarchical processing in the visual cortex [20]. In this formulation, synaptic weights are updated after inference to minimize the prediction error of the model. The connection with variational inference was established later, connecting PC to hierarchical Gaussian generative models [36, 37]. As a result of these independent but deeply related lines of work, a general theory has unified these views by showing that iterative updating schemes can be described as a process of minimizing a variational free energy [38], also known in machine learning as the evidence lower bound (ELBO) [39]. A sketch of the timeline of main breakthroughs in PC history is given in Fig. 2.
@@ 2.1 What Defines Predictive Coding?
The PC literature has historically focused on a specific class of models that make two key simplifying assumptions: first, it typically assumes that all probability distributions are multivariate Gaussian; second, it often employs a hierarchical structure where latent variables follow a first-order Markov assumption—each layer depends only on the layer immediately above it. These assumptions, while restrictive, provide significant computational advantages: Gaussian distributions are conjugate to themselves under linear transformations, enabling closed-form updates, and the Markov structure allows for local computations between adjacent layers.
Let $\mathbf{o}=g(\mathbf{x},\mathbf{w})$ be a generative model, where a vector of latent variables $\mathbf{x}$ and a set of parameters $\mathbf{w}$ are used to generate an observation $\mathbf{o}$. In our case, we consider generative models with $L$ layers, whose joint probability follows the hierarchical dependencies:
@@
We further consider each conditional probability distribution $p(\mathbf{x}_{\ell}\mid\mathbf{x}_{\ell+1})$ to be a multivariate Gaussian distribution, whose mean is given by a transformation $g_{\ell}$ of the latent variables of the level above. In the majority of the literature, the map $g_{\ell}$ is a composition of an activation function (such as ReLU), and a transformation matrix $\mathbf{w}_{\ell}$ (with bias $\mathbf{b}_{\ell}$ possibly folded into this matrix), which results in a linear map. As a result, we then, more formally, arrive at the following:
$p(\mathbf{x}_{L})=\mathcal{N}(\mathbf{x}_{L};\boldsymbol{\mu}_{L},\boldsymbol{\Sigma}_{L}),\quad p(\mathbf{x}_{\ell}\mid\mathbf{x}_{\ell+1})=\mathcal{N}(\mathbf{x}_{\ell};g_{\ell+1}(\mathbf{x}_{\ell+1}),\boldsymbol{\Sigma}_{\ell}).$ (2)
@@
Given the above, we now have the structure of the main object of study that centrally characterizes the PC literature, that is, a hierarchical generative model, which lives in a continuous state space and whose probability distributions are Gaussian in nature. However, what also defines PC is the process that is used to invert the generative model, which means estimating $p(\mathbf{x}_{1},\dots,\mathbf{x}_{L}\mid\mathbf{o})$ via an approximate posterior $q(\mathbf{x}_{1},\dots,\mathbf{x}_{L})$. To do this, we need to appeal to two different approximation results: first, we leverage a mean-field approximation, which allows us to assume that the variational posterior factorizes into conditionally independent posteriors $q(\mathbf{x}_{\ell})$; then, through the Laplace approximation, we assume that approximate posterior distributions are Gaussian in form [43]. At this point, the model can be inverted by minimizing the resulting variational free energy via gradient descent or fixed-point iterations.
As the resulting variational free energy has a quadratic form [44, 41], its gradients correspond to the linear and weighted prediction errors defined in the original computational model by Rao and Ballard [20], which was notably developed before PC was cast as variational learning and inference [41]. We have now introduced all of the core concepts needed to provide a definition of PC. This definition could be used as an umbrella for all of the variations on PC in the literature.
@@
    it maximizes the model evidence $\log p(\mathbf{o})$ by minimizing a variational free energy,
@@
    the posterior distributions of the nodes of the hierarchical structure are factorized via a mean-field approximation, and
@@
Note that the above definition does not say anything explicitly about prediction error or properties such as locality, which, as mentioned earlier, are commonly used to describe PC. This is because they are not foundational to PC but rather consequences of the commitment to the aforementioned generative model: the mean field approximation enforces independence, and hence, results in locality in the update rules; the Laplace approximation simplifies variational free energy to a quadratic function, meaning that its gradients are linear prediction errors.
It is crucial to note that the above definition of PC does not mean, however, that the algorithms surveyed here are optimal and cannot be improved: the above definition is quite general and does not impose any constraint on the exact computation of the posteriors as well as the optimization technique(s) used to minimize the variational free energy. It also does not mean that such conditions cannot be relaxed: recent works have also explored alternatives to Gaussian and Markov assumptions. Common alternatives use amortized computations using a similar technique to VAEs [45, 46, 47, 48], rely on different probability distributions [46, 49] or use models with different architectures [25, 50]. It is also worth noting a recent trend that relies on sampling algorithms to improve the generative capacities of PC models [51, 52, 47]. However, we believe the classical framework remains essential for understanding the foundational principles that give PC its characteristic properties.
3 Implementations of Predictive Coding